OpenStudy (anonymous):

solve using the multiplication principle -1/2x = -5/6

OpenStudy (anonymous):

DO CROSS MULTIPLICATION HERE:

OpenStudy (anonymous):

can you show the work

OpenStudy (anonymous):

multiply both sides by -2

jhonyy9 (jhonyy9):

multiply bot sides by (-6x) and you get ((-1)(-6x))/2x) = ((-5)(-6x))/6 --- (6x/2x) = (30x/6) --- after you simplify you get 3=5x --- x=3/5